# Low Regularity of Self-Similar Solutions of Two-Dimensional Riemann problems with Shocks for the Isentropic Euler system

### v1

@@ Abstract.
We are concerned with the low regularity of self-similar solutions of two-dimensional Riemann problems for the isentropic Euler system. We establish a general framework for the analysis of the local regularity of such solutions for a class of two-dimensional Riemann problems for the isentropic Euler system, which includes the regular shock reflection problem, the Prandtl reflection problem, the Lighthill diffraction problem, and the four-shock Riemann problem. We prove that it is not possible that both the density and the velocity are in $H^{1}$ in the subsonic domain for the self-similar solutions of these problems in general. This indicates that the self-similar solutions of the Riemann problems with shocks for the isentropic Euler system are of much more complicated structure than those for the Euler system for potential flow; in particular, the density and the velocity are not necessarily continuous in the subsonic domain. The proof is based on a regularization of the isentropic Euler system to derive the transport equation for the vorticity, a renormalization argument extended to the case of domains with boundary, and DiPerna-Lions-type commutator estimates.
@@
Low regularity, transonic shock, free boundary, compressible flow, potential flow, self-similar, conservation laws, shock reflection-diffraction, regular shock reflection, Prandtl reflection, Lighthill diffraction, nonlinear, fine properties
@@ 1. Introduction
More recently, the existence of global regular reflection solutions was established for the Euler system for potential flow in [14, 15], and further properties of these solutions were proved in [1, 17]. In particular, the solution has a high regularity in the subsonic domain $\Omega$ (where the solution is not a constant state); see Figs. 4.24.2 in which density $\rho$ and velocity ${\mathbf{v}}$ are in $C^{\alpha}(\overline{\Omega})\cap C^{\infty}(\Omega)$ for some $\alpha\in(0,1)$. More precisely, solution $(\rho,{\mathbf{v}})$ is in a weighted Hölder space which implies that $(\rho,{\mathbf{v}})$ is in the Sobolev space $W^{1,p}(\Omega)$ for some $p>2$. In addition, the regularity of the curved reflected-diffracted shock was shown to be $C^{\infty}$ in the interior and $C^{2,\alpha}$ up to the endpoints, which is also expected to hold for the case of the isentropic Euler system. See also [2, 3, 26] for the Prandtl reflection-diffraction configuration, [13] for the Lighthill diffraction problem, and [12] for the four-shock Riemann problem.
@@
In this paper, under the natural assumptions on the regularity of the self-similar solutions in the subsonic domain $\Omega$ near the shocks, we rigorously prove that it is not possible indeed that both the density and the velocity are in $H^{1}(\Omega)$. The argument is based on the vortical singularity calculation in [44]. We first apply this calculation to the regularized solutions carefully for which the calculation can be rigorously justified; however, the additional error terms appear due to the regularization. Then we develop DiPerna-Lions-type commutator estimates to control the error terms when the regularization parameter tends to zero. With this, we then prove the lower regularity property of the self-similar solutions, by employing renormalization argument. These self-similar solutions include the above-mentioned shock reflection problems: the regular shock reflection problem, the Prandtl reflection problem, the Lighthill diffraction problem, and the four-shock Riemann problem.
This indicates that the self-similar solutions of the Riemann problems with shocks for the isentropic Euler system have much more complicated structure than the corresponding solutions for the Euler system for potential flow; in particular, the density and the velocity are not necessarily continuous in $\Omega$, at least their continuity can not be obtained directly by the Sobolev embeddings. On the other hand, this argument allows the possibility that $(\rho,{\mathbf{v}})\in W^{1,p}(\Omega)$ for some or even all $p\in[1,2)$, in which case there are no shocks in $\Omega$.
@@ 2.1. The Isentropic Euler System
where $h(\rho)=\frac{\rho^{\gamma-1}}{\gamma-1}$ is the enthalpy. Note that the speed of sound is $c=\rho^{\frac{\gamma-1}{2}}=\sqrt{(\gamma-1)h(\rho)}$.
@@ 2.2. Entropy Solutions and the Rankine-Hugoniot Conditions
Suppose that $S$ is a smooth curve in domain $\Lambda$. An entropy solution of (2.3)–(2.4), which is $C^{1}$ near and up to $S$ on both sides of $S$, satisfies the following Rankine-Hugoniot (R-H) conditions on $S$:
$[\rho{\mathbf{v}}\cdot\boldsymbol{\nu}]=0,\quad\,[(\rho{\mathbf{v}}\cdot\boldsymbol{\nu}){\mathbf{v}}+p\boldsymbol{\nu}]={\bf 0}\qquad\,\,\mbox{on $S$},$ (2.9)
[…]
If $\rho^{+}{\mathbf{v}}^{+}\cdot\boldsymbol{\nu}=0$ on $S$, then $\rho^{-}{\mathbf{v}}^{-}\cdot\boldsymbol{\nu}=0$ on $S$ by the first equality in (2.9), and $[p]=0$ from the second equality in (2.9) which implies that $[\rho]=0$. This discontinuity is called a vortex sheet. In this case, $[{\mathbf{v}}\cdot\boldsymbol{\tau}]\neq 0$ on $S$, unless the solution is continuous across $S$, where $\boldsymbol{\tau}$ is a unit tangent vector to $S$.
If $\rho^{+}{\mathbf{v}}^{+}\cdot\boldsymbol{\nu}\neq 0$ on $S$, then $\rho^{-}{\mathbf{v}}^{-}\cdot\boldsymbol{\nu}\neq 0$ on $S$ by the first equality in (2.9). This discontinuity is called a shock. In this case, $\rho{\mathbf{v}}\cdot\boldsymbol{\nu}$ is continuous across $S$ from the first equality in (2.9), and $[{\mathbf{v}}\cdot\boldsymbol{\tau}]=0$ and $[p]=-\rho{\mathbf{v}}\cdot\boldsymbol{\nu}[{\mathbf{v}}\cdot\boldsymbol{\nu}]$ from the second equality in (2.9). This implies that $[{\mathbf{v}}\cdot\boldsymbol{\nu}]\neq 0$, unless the solution is continuous across $S$, so that $[\rho]\neq 0$ from the first equality in (2.9). Also, it follows from $\rho^{\pm}>0$ on $S$ that $({\mathbf{v}}^{+}\cdot\boldsymbol{\nu})({\mathbf{v}}^{-}\cdot\boldsymbol{\nu})>0$ on $S$.
@@
[…]
@@
are satisfied for any test function $\zeta\in C^{\infty}_{\rm c}(\mathbb{R}^{2};\,\mathbb{R}^{2})$. Note that here the test function is not required to vanish on $\partial\Lambda$. Based on that, we define the notion of entropy solutions of the boundary value problem (2.3)–(2.4) and (2.13) as follows:
@@
Figure 3.1. The Riemann Problem in a General Setting
More precisely, as shown in Fig. 3.1, we consider a Riemann problem in a self-similar wedge domain $\Lambda$ in $\mathbb{R}^{2}$ with the wedge-vertex being the origin, which also includes the case of the whole space. That is, in polar coordinates,
either Λ=2 or $\mbox{either $\,\,\Lambda=\mathbb{R}^{2}\,\,\,$ or $\,\,\Lambda=\big{\{}(r,\theta)\;:\;\theta^{-}<\theta<\theta^{+}\big{\}}$}.$ (3.1)
For a Riemann problem in a domain $\Lambda$ whose boundary contains a wedge-boundary, the initial data are the given constant velocity and density in each sub-sector of $\Lambda$ such that the Rankine-Hugoniot conditions hold on the lines separating them. Motivated by the expected configurations of self-similar solutions of the Riemann problems in §4, we consider solutions in self-similar coordinates of the following form, in terms of the density and pseudo-velocity $(\rho,{\mathbf{v}})$ with $\mathcal{N}_{r}(\Gamma):=\{{\boldsymbol{\xi}}\,:\,\mbox{\rm dist}\{{\boldsymbol{\xi}},\Gamma\}<r\}$:
@@
We say that the vector function $(\rho,{\mathbf{v}})$ on $\Lambda$ is an entropy solution of a Riemann problem of admissible structure if $(\rho,{\mathbf{v}})\in L^{\infty}(\Lambda)$, with $\rho\in BV_{\rm loc}(\Lambda\cap{\mathcal{N}}_{r}(\partial\Lambda))$ for some $r>0$, is an entropy solution of system (2.3)–(2.4) with the slip boundary condition (2.13) in the sense of Definition 2.2, which satisfies the following properties:
@@
    There exists an open, bounded, connected set $\Omega\subset\Lambda$, an integer $M\geq 1$, and open connected sets $\Lambda_{i}$, $i=1,\cdots,M$, pairwise disjoint, such that
    $\overline{\Lambda}\setminus\Omega=\cup_{i=1}^{M}\overline{\Lambda_{i}},$
    and $(\rho,{\mathbf{v}})$ is a constant state in each $\Lambda_{i}$, i.e., $(\rho,{\mathbf{v}})({\boldsymbol{\xi}})=(\rho_{i},{\mathbf{u}}_{i}-{\boldsymbol{\xi}})$ in $\Lambda_{i}$, where $\rho_{i}>0$ is a constant and ${\mathbf{u}}_{i}$ is a constant vector.
@@
    If, for $i\neq j$, sets $\Lambda_{i}$ and $\Lambda_{j}$ have common boundary within $\Lambda$; that is, if $\partial\Lambda_{i}\cap\partial\Lambda_{j}\cap\Lambda\neq\emptyset$, then the corresponding constant states are not equal to each other: $(\rho_{i},{\mathbf{v}}_{i})\neq(\rho_{j},{\mathbf{v}}_{j})$.
@@
    $\partial\Omega$ is Lipschitz. Denote
    $\Gamma^{\mbox{\rm\footnotesize ext}}=\partial\Omega\cap\partial\Lambda,\qquad\Gamma^{\mbox{\rm\footnotesize int}}=\partial\Omega\cap\Lambda,$
    so that $\partial\Omega=\Gamma^{\mbox{\rm\footnotesize ext}}\cup\Gamma^{\mbox{\rm\footnotesize int}}$. It is possible that $\Gamma^{\mbox{\rm\footnotesize ext}}=\emptyset$ (in particular, this is true when $\partial\Lambda=\emptyset$).
@@
    If $\partial\Lambda\neq\emptyset$, then ${\Gamma^{\mbox{\rm\footnotesize ext}}}=\cup_{i=1}^{N_{1}}\overline{\Gamma^{\mbox{\rm\footnotesize ext}}_{i}}$, where ${N_{1}}\geq 1$ and each $\Gamma^{\mbox{\rm\footnotesize ext}}_{i}$ is a relatively open segment of straight line. Segments $\{\Gamma^{\mbox{\rm\footnotesize ext}}_{i}\}_{i=1}^{N_{1}}$ are disjoint and, if $\Gamma^{\mbox{\rm\footnotesize ext}}_{i}$ and $\Gamma^{\mbox{\rm\footnotesize ext}}_{j}$ have a common endpoint, then the interior angle (for $\Omega$) is within $(0,2\pi)$ at that point. Since $\Lambda$ is a wedge domain with the wedge-vertex at the origin, only one pair among segments $\{\Gamma^{\mbox{\rm\footnotesize ext}}_{i}\}$ may have a common endpoint that is $\mathcal{P}_{0}=(0,0)$.
@@
    $\overline{\Gamma^{\mbox{\rm\footnotesize int}}}=\cup_{i=1}^{N_{2}}\overline{\Gamma^{\mbox{\rm\footnotesize int}}_{i}}$, where ${N_{2}}\geq 1$ and each $\Gamma^{\mbox{\rm\footnotesize int}}_{i}$ is a relatively open segment of curve, and segments $\{\Gamma^{\mbox{\rm\footnotesize int}}_{i}\}_{i=1}^{N_{2}}$ are disjoint. Each $\Gamma^{\mbox{\rm\footnotesize int}}_{i}$ is $C^{2}$ in its relative interior and $C^{1}$ up to the endpoints. Moreover, $\mathcal{P}_{0}\notin\overline{\Gamma^{\mbox{\rm\footnotesize int}}}$.
@@
    There exists a point $\hat{P}$ in the relative interior of $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ such that the curvature of $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ is non-zero at $\hat{P}$, and $({\mathbf{v}}\cdot{\boldsymbol{\tau}})(\hat{P})\neq 0$, where ${\boldsymbol{\tau}}(\hat{P})$ is a unit tangent vector to $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ at $\hat{P}$.
@@
Let $(\rho,{\mathbf{v}})$ be a solution of a Riemann problem of admissible structure in the sense of Definition 3.1. Assume that $(\rho,{\mathbf{v}})$ satisfy
@@
Then it is not possible that $(\rho,{\mathbf{v}})\in H^{1}(\Omega)$.
@@
We divide the proof into six steps.
1. We prove the theorem by establishing that, under the assumption:
@@
the calculations of vortical singularity can rigorously be justified, which leads to the contradiction.
Under assumption (3.2), vorticity $\omega=\nabla\times{\mathbf{v}}$ satisfies $\omega\in L^{2}(\Omega)$ and
@@
for any $f\in C^{1}(\mathbb{R})$ with $g$ defined by (3.8). From the weak form (2.6) of the conservation law of mass, which holds for all $\phi$ specified in Definition 2.2, and the regularity assumption in (3.2), we obtain that $\rho{\mathbf{v}}\cdot{\boldsymbol{\nu}}=0$ holds ${\mathcal{H}}^{1}$a.e. on $\Gamma^{\mbox{\rm\footnotesize ext}}$ in the sense of traces. Then, using assumption (ii), we recover condition (2.13) on $\partial\Lambda\cap\partial\Omega$:
@@
for any $f\in C^{1}(\mathbb{R}^{1})$ with $\|f^{\prime}\|_{C^{0,1}(\mathbb{R}^{1})}<\infty$.
@@
$\rho f(X){\mathbf{v}}\cdot{\boldsymbol{\nu}}=0\qquad\mbox{on $\Gamma^{\mbox{\rm\footnotesize ext}}$ in the weak sense.}$.
@@
$\Gamma^{\mbox{\rm\footnotesize ext}}_{i}\cap B_{(L+1)r_{1}}(\mathcal{P}_{0})=\emptyset\qquad\mbox{if $\mathcal{P}_{0}$ is not an endpoint of }\Gamma^{\mbox{\rm\footnotesize ext}}_{i}.$ (3.13)
@@
Using (3.2) and assumption (ii), and noting that ${\mathbf{v}}\cdot{\boldsymbol{\nu}}=0$ to make the odd extension of ${\mathbf{v}}\cdot{\boldsymbol{\nu}}$ across $\Gamma^{\mbox{\rm\footnotesize ext}}_{i}$, we find that the extended $(\rho,{\mathbf{v}})$ satisfies $(\rho,{\mathbf{v}})\in H^{1}((\Omega\cup\overline{V_{i}^{r_{1}}})^{0})\cap L^{\infty}(\Omega\cup\overline{V_{i}^{r_{1}}})$ and $\rho\geq C_{0}^{-1}$ on $\Omega\cup\overline{V_{i}^{r_{1}}}$.
@@
Under assumptions (i)–(ii) of Theorem 3.2 and (3.2), there exists $r_{1}>0$ depending only on $\Omega$ such that, for any $r\in(0,r_{1}]$, there exists an extension of $(\rho,{\mathbf{v}})$ into $\Omega^{\rm ext}_{r}$, still denoted as $(\rho,{\mathbf{v}})$, such that
@@
${\mathbf{v}}_{\varepsilon}={\mathbf{v}}*\eta_{\varepsilon},\quad\omega_{\varepsilon}=\omega*\eta_{\varepsilon},\quad\rho_{\varepsilon}=\rho*\eta_{\varepsilon}\quad(i(\rho))_{\varepsilon}=i(\rho)*\eta_{\varepsilon}\qquad\,\,\mbox{for $\varepsilon\in(0,\frac{\delta}{2})$}.$
@@
$\rho_{\varepsilon},\,{\mathbf{v}}_{\varepsilon},\,(i(\rho))_{\varepsilon}\in H^{1}(\Omega^{r})\cap L^{\infty}(\Omega^{r}),\quad\|(\rho_{\varepsilon}\,{\mathbf{v}}_{\varepsilon})\|_{L^{\infty}(\Omega^{r})}\leq\|(\rho\,{\mathbf{v}})\|_{L^{\infty}(\Omega^{\rm ext}_{r})},$ (3.23)

### v2

@@ Abstract.
We are concerned with the low regularity of self-similar solutions of two-dimensional Riemann problems for the isentropic Euler system. We establish a general framework for the analysis of the local regularity of such solutions for a class of two-dimensional Riemann problems for the isentropic Euler system, which includes the regular shock reflection problem, the Prandtl reflection problem, the Lighthill diffraction problem, and the four-shock Riemann problem. We prove that the velocity is not in $H^{1}$ in the subsonic domain for the self-similar solutions of these problems in general. This indicates that the self-similar solutions of the Riemann problems with shocks for the isentropic Euler system are of much more complicated structure than those for the Euler system for potential flow; in particular, the velocity is not necessarily continuous in the subsonic domain. The proof is based on a regularization of the isentropic Euler system to derive the transport equation for the vorticity, a renormalization argument extended to the case of domains with boundary, and DiPerna-Lions-type commutator estimates.
@@
Low regularity, transonic shock, free boundary, compressible flow, potential flow, self-similar, conservation laws, shock reflection-diffraction, regular shock reflection, Prandtl reflection, Lighthill diffraction, nonlinear, fine properties.
*Corresponding author: Gui-Qiang G. Chen
@@ 1. Introduction
More recently, the existence of global regular reflection solutions was established for the Euler system for potential flow in [14, 15], and further properties of these solutions were proved in [1, 17]. In particular, the solution has a high regularity in the subsonic domain $\Omega$ (where the solution is not a constant state); see Figs. 33 in which density $\rho$ and velocity $\mathbf{v}$} are in $C^{\alpha}(\overline{\Omega})\cap C^{\infty}(\Omega)$ for some $\alpha\in(0,1)$. More precisely, solution $(\rho,{\mathbf{v}})$ is in a weighted Hölder space which implies that $(\rho,{\mathbf{v}})$ is in the Sobolev space $W^{1,p}(\Omega)$ for some $p>2$. In addition, the regularity of the curved reflected-diffracted shock was shown to be $C^{\infty}$ in the interior and $C^{2,\alpha}$ up to the endpoints, which is also expected to hold for the case of the isentropic Euler system. See also [2, 3, 26] for the Prandtl reflection-diffraction configuration, [13] for the Lighthill diffraction problem, and [12] for the four-shock Riemann problem.
@@
In this paper, under the natural assumptions on the regularity of the self-similar solutions in the subsonic domain $\Omega$ near the shocks, we rigorously prove that the velocity is not in $H^{1}(\Omega)$ indeed. The argument is based on the vortical singularity calculation in [44]. We first apply this calculation to the regularized solutions carefully for which the calculation can be rigorously justified; however, the additional error terms appear due to the regularization. Then we develop DiPerna-Lions-type commutator estimates to control the error terms when the regularization parameter tends to zero. With this, we then prove the lower regularity property of the self-similar solutions, by employing renormalization argument. These self-similar solutions include the above-mentioned shock reflection problems: the regular shock reflection problem, the Prandtl reflection problem, the Lighthill diffraction problem, and the four-shock Riemann problem.
This indicates that the self-similar solutions of the Riemann problems with shocks for the isentropic Euler system have much more complicated structure than the corresponding solutions for the Euler system for potential flow; in particular, the velocity is not necessarily continuous in $\Omega$, at least their continuity can not be obtained directly by the Sobolev embeddings. On the other hand, this argument allows the possibility that $(\rho,{\mathbf{v}})\in W^{1,p}(\Omega)$ for some or even all $p\in[1,2)$, in which case there are no shocks in $\Omega$.
@@ 2.1. The Isentropic Euler System
where $h(\rho)=\frac{\rho^{\gamma-1}-1}{\gamma-1}$ is the enthalpy. Note that the speed of sound is $c=\rho^{\frac{\gamma-1}{2}}=\sqrt{(\gamma-1)h(\rho)+1}$.
@@ 2.2. Entropy Solutions and the Rankine-Hugoniot Conditions
Suppose that $S$ is a smooth curve in domain $\Lambda$. An entropy solution of (2.3)–(2.4), which is $C^{1}$ in the open domains near $S$ and $C^{0}$ up to $S$ on both sides, satisfies the following Rankine-Hugoniot (R-H) conditions on $S$:
$\rho{\mathbf{v}}\cdot\boldsymbol{\nu}]=0,\quad\,[(\rho{\mathbf{v}}\cdot\boldsymbol{\nu}){\mathbf{v}}+p\boldsymbol{\nu}]={\bf 0}\qquad\,\,\mbox{on $S$},$ (2.9)
where $\boldsymbol{\nu}$ is a unit normal to $S$, and $\cdot\,$] denotes the difference of the concerned quantity across $S$. Denote by $(\rho^{\pm},{\mathbf{v}}^{\pm})$ the values of solution $(\rho,{\mathbf{v}})$ on the $\pm$ sides of $S$. Assume that $\rho^{\pm}>0$ on $S$ and some of $(\rho,{\mathbf{v}})$ are discontinuous across $S$.
If $\rho^{+}{\mathbf{v}}^{+}\cdot\boldsymbol{\nu}=0$ on $S$, then $\rho^{-}{\mathbf{v}}^{-}\cdot\boldsymbol{\nu}=0$ on $S$ by the first equality in (2.9), and $[p]=0$ from the second equality in (2.9) which implies that $\rho]=0$. This discontinuity is called a vortex sheet. In this case, ${\mathbf{v}}\cdot\boldsymbol{\tau}]\neq 0$ on $S$, unless the solution is continuous across $S$, where $\boldsymbol{\tau}$ is a unit tangent vector to $S$.
If $\rho^{+}{\mathbf{v}}^{+}\cdot\boldsymbol{\nu}\neq 0$ on $S$, then $\rho^{-}{\mathbf{v}}^{-}\cdot\boldsymbol{\nu}\neq 0$ on $S$ by the first equality in (2.9). This discontinuity is called a shock. In this case, $\rho{\mathbf{v}}\cdot\boldsymbol{\nu}$ is continuous across $S$ from the first equality in (2.9), and ${\mathbf{v}}\cdot\boldsymbol{\tau}]=0$ and $p]=-\rho{\mathbf{v}}\cdot\boldsymbol{\nu}[{\mathbf{v}}\cdot\boldsymbol{\nu}]$ from the second equality in (2.9). This implies that ${\mathbf{v}}\cdot\boldsymbol{\nu}]\neq 0$, unless the solution is continuous across $S$, so that $\rho]\neq 0$ from the first equality in (2.9). Also, it follows from $\rho^{\pm}>0$ on $S$ that $({\mathbf{v}}^{+}\cdot\boldsymbol{\nu})({\mathbf{v}}^{-}\cdot\boldsymbol{\nu})>0$ on $S$.
@@
${\mathbf{v}}\cdot{\boldsymbol{\nu}}=0\qquad\mbox{on $\partial\Lambda$},$ (2.13)
@@
are satisfied for any test function $\zeta\in C^{\infty}_{\rm c}(\mathbb{R}^{2};\,\mathbb{R}^{2})$. Note that here the test function is not required to vanish on $\partial\Lambda$.
From now on, we define the $r$-neighborhood of a curve $\Gamma$ as
$\mathcal{N}_{r}(\Gamma):=\{{\boldsymbol{\xi}}\,:\,\mbox{\rm dist}\{{\boldsymbol{\xi}},\Gamma\}<r\}.$ (2.15)
Based on that, we define the notion of entropy solutions of the boundary value problem (2.3)–(2.4) and (2.13) as follows:
@@
Figure 1. The Riemann Problem in a General Setting
More precisely, as shown in Fig. 1, we consider a Riemann problem in a self-similar wedge domain $\Lambda$ in $\mathbb{R}^{2}$ with the wedge-vertex being the origin, which also includes the case of the whole space. That is, in polar coordinates,
$\mbox{either $\,\,\Lambda=\mathbb{R}^{2}\,\,\,$ or $\,\,\Lambda=\big\{(r,\theta)\;:\;\theta^{-}<\theta<\theta^{+}\big\}$}.$ (3.1)
For a Riemann problem in a domain $\Lambda$ whose boundary contains a wedge-boundary, the initial data are the given constant velocity and density in each sub-sector of $\Lambda$ such that the Rankine-Hugoniot conditions hold on the lines separating them. Motivated by the expected configurations of self-similar solutions of the Riemann problems in §4, we consider solutions in self-similar coordinates of the following form, in terms of the density and pseudo-velocity $(\rho,{\mathbf{v}})$:
@@
We say that the vector function $(\rho,{\mathbf{v}})$ on $\Lambda$ is an entropy solution of the Riemann problem of admissible structure if $(\rho,{\mathbf{v}})\in L^{\infty}(\Lambda)$, with $\rho\in BV_{\rm loc}(\Lambda\cap{\mathcal{N}}_{r}(\partial\Lambda))$ for some $r>0$, is an entropy solution of system (2.3)–(2.4) with the slip boundary condition (2.13) in the sense of Definition 2.2, which satisfies the following properties:
@@
    There exist an open, bounded, connected set $\Omega\subset\Lambda$, an integer $M\geq 1$, and open connected sets $\Lambda_{i}$, $i=1,\cdots,M$, pairwise disjoint, such that ΛΩ=_i=1^MΛ_i, and $(\rho,{\mathbf{v}})$ is a constant state in each $\Lambda_{i}$, i.e., $(\rho,{\mathbf{v}})({\boldsymbol{\xi}})=(\rho_{i},{\mathbf{u}}_{i}-{\boldsymbol{\xi}})$ in $\Lambda_{i}$, where $\rho_{i}>0$ is a constant and ${\mathbf{u}}_{i}$ is a constant vector.
@@
    If, for $i\neq j$, sets $\Lambda_{i}$ and $\Lambda_{j}$ have common boundary within $\Lambda$, i.e., if $\partial\Lambda_{i}\cap\partial\Lambda_{j}\cap\Lambda\neq\emptyset$, then the corresponding constant states are not equal to each other: $(\rho_{i},{\mathbf{v}}_{i})\neq(\rho_{j},{\mathbf{v}}_{j})$.
@@
    $\partial\Omega$ is Lipschitz. Denote Γ^ext=Ω∩∂Λ,   Γ^int=ΩΛ, so that $\partial\Omega=\Gamma^{\mbox{\rm\footnotesize ext}}\cup\Gamma^{\mbox{\rm\footnotesize int}}$. It is possible that $\Gamma^{\mbox{\rm\footnotesize ext}}=\emptyset$ (in particular, this is true when $\partial\Lambda=\emptyset$).
@@
    If $\partial\Lambda\neq\emptyset$, then ${\Gamma^{\mbox{\rm\footnotesize ext}}}=\cup_{i=1}^{N_{1}}\overline{\Gamma^{\mbox{\rm\footnotesize ext}}_{i}}$, where ${N_{1}}\geq 1$, each $\Gamma^{\mbox{\rm\footnotesize ext}}_{i}$ is a relatively open segment of straight line. Segments $\{\Gamma^{\mbox{\rm\footnotesize ext}}_{i}\}_{i=1}^{N_{1}}$ are disjoint and, if $\Gamma^{\mbox{\rm\footnotesize ext}}_{i}$ and $\Gamma^{\mbox{\rm\footnotesize ext}}_{j}$ have a common endpoint, then the interior angle (for $\Omega$) is within $(0,2\pi)$ at that point. Since $\Lambda$ is a wedge domain with the wedge-vertex at the origin, only one pair among segments $\{\Gamma^{\mbox{\rm\footnotesize ext}}_{i}\}$ may have a common endpoint that is $\mathcal{P}_{0}=(0,0)$.
@@
    $\overline{\Gamma^{\mbox{\rm\footnotesize int}}}=\cup_{i=1}^{N_{2}}\overline{\Gamma^{\mbox{\rm\footnotesize int}}_{i}}$, where ${N_{2}}\geq 1$, each $\Gamma^{\mbox{\rm\footnotesize int}}_{i}$ is a relatively open segment of curve, and segments $\{\Gamma^{\mbox{\rm\footnotesize int}}_{i}\}_{i=1}^{N_{2}}$ are disjoint. Each $\Gamma^{\mbox{\rm\footnotesize int}}_{i}$ is $C^{2}$ in its relative interior and $C^{1}$ up to the endpoints. Moreover, $\mathcal{P}_{0}\notin\overline{\Gamma^{\mbox{\rm\footnotesize int}}}$.
@@
[…]
@@
Let $(\rho,{\mathbf{v}})$ be a solution of the Riemann problem of admissible structure in the sense of Definition 3.1. Assume that $(\rho,{\mathbf{v}})$ satisfy
@@
Then ${\mathbf{v}}\notin H^{1}(\Omega)$.
@@
We divide the proof into seven steps. In the first six steps, we prove that both $\rho$ and $\mathbf{v}$} can not be in $H^{1}(\Omega)$:
$(\rho,{\mathbf{v}})\notin H^{1}(\Omega).$ (3.2)
This will be proved by showing that, under the assumption that
@@
the calculations of vortical singularity can rigorously be justified, which leads to the contradiction. Then, in Step 7, we prove that $\mathbf{v}$} is not in $H^{1}(\Omega)$ indeed: ${\mathbf{v}}\notin H^{1}(\Omega)$.
1. In Steps 1–6 of this proof, we assume that (3.3) holds. Then vorticity $\omega=\nabla\times{\mathbf{v}}$ satisfies $\omega\in L^{2}(\Omega)$ and
@@
for any $f\in C^{1}(\mathbb{R})$ with $g$ defined by (3.9). From the weak form (2.6) of the conservation law of mass, which holds for all $\phi$ specified in Definition 2.2, and the regularity assumption (3.3), we obtain that $\rho{\mathbf{v}}\cdot{\boldsymbol{\nu}}=0$ holds ${\mathcal{H}}^{1}$a.e. on $\Gamma^{\mbox{\rm\footnotesize ext}}$ in the sense of traces. Then, using assumption (ii), we recover condition (2.13) on $\partial\Lambda\cap\partial\Omega$:
@@
for any $f\in C^{1}(\mathbb{R})$ with $\|f^{\prime}\|_{C^{0,1}(\mathbb{R})}<\infty$.
@@
$\rho f(X){\mathbf{v}}\cdot{\boldsymbol{\nu}}=0\qquad\mbox{on $\Gamma^{\mbox{\rm\footnotesize ext}}$ in the weak sense.}$
@@
[…]
@@
Using (3.3) and (ii), and noting that ${\mathbf{v}}\cdot{\boldsymbol{\nu}}=0$ to make the odd extension of ${\mathbf{v}}\cdot{\boldsymbol{\nu}}$ across $\Gamma^{\mbox{\rm\footnotesize ext}}_{i}$, we find that the extended $(\rho,{\mathbf{v}})$ satisfies $(\rho,{\mathbf{v}})\in H^{1}((\Omega\cup\overline{V_{i}^{r_{1}}})^{0})\cap L^{\infty}(\Omega\cup\overline{V_{i}^{r_{1}}})$ and $\rho\geq C_{0}^{-1}$ on $\Omega\cup\overline{V_{i}^{r_{1}}}$.
@@
Under assumptions (i)–(ii) of Theorem 3.2 and (3.3), there exists $r_{1}>0$ depending only on $\Omega$ such that, for any $r\in(0,r_{1}]$, there is an extension of $(\rho,{\mathbf{v}})$ into $\Omega^{\rm ext}_{r}$, still denoted as $(\rho,{\mathbf{v}})$, so that
@@
${\mathbf{v}}_{\varepsilon}={\mathbf{v}}*\eta_{\varepsilon},\quad\omega_{\varepsilon}=\omega*\eta_{\varepsilon},\quad\rho_{\varepsilon}=\rho*\eta_{\varepsilon}\quad(h(\rho))_{\varepsilon}=h(\rho)*\eta_{\varepsilon}\qquad\,\,\mbox{for $\varepsilon\in(0,\frac{\delta}{2})$}.$
@@
$\rho_{\varepsilon},\,{\mathbf{v}}_{\varepsilon},\,(h(\rho))_{\varepsilon}\in H^{1}(\Omega^{r})\cap L^{\infty}(\Omega^{r}),\quad\|(\rho_{\varepsilon},\,{\mathbf{v}}_{\varepsilon})\|_{L^{\infty}(\Omega^{r})}\leq\|(\rho,\,{\mathbf{v}})\|_{L^{\infty}(\Omega^{\rm ext}_{r})},$ (3.24)